OpenStudy (anonymous):

Factor: 9x2 - 144x + 64

OpenStudy (anonymous):

use the formula (a-b)^2=a^2-2ab+b^2

OpenStudy (anonymous):

find out the value of a and b from your expression

OpenStudy (maheshmeghwal9):

u have to factorize it as (a+x)(b+y)=0

OpenStudy (maheshmeghwal9):

\[[9x^2-144x+63]+1\] now factorize

OpenStudy (maheshmeghwal9):

\[3[3x^2-48x+21]+1\]

OpenStudy (unklerhaukus):

or just use the quadratic formula\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

out of these... (3x + 4)(3x - 8) (3x + 8)2 (3x - 8)2 prime

OpenStudy (anonymous):